Consider a normal population with a mean of 50 and standard deviation of 2.Find the probability that a randomly picked number is between 49.1 and 50.15.
Answer by Boreal(15235)

z=(x-mean)/sd
so z is between (49.1-50)/2=-.9/2=-0.45
and (50.15-50)/2=0.15/2=+0.075
that probability is 0.2035